OpenStudy (anonymous):

If 4^x=7 then 4^-2x =?

OpenStudy (anonymous):

square both sides by \[2\] then invert

OpenStudy (anonymous):

since \(\large 4^x=7 \rightarrow x=log_47 \) then \(\large 4^{-2x}=4^{-2 \cdot (log_47)} \) simplify....

OpenStudy (anonymous):

\[(4^x)^2=7^2\] \[\frac{ 1 }{ (4^x)^2 }=4^{-2x}\] so \[\frac{ 1 }{ 7^2 }\]

OpenStudy (anonymous):

if i square both sides ill get 16x^2=49

OpenStudy (anonymous):

then apply inversion rule \[a^{-1}=\frac{ 1 }{ a }\]

OpenStudy (anonymous):

\[\frac{ 1 }{ 16x^{-2} }=\frac{ 1 }{ 49 }\]